4. At the corners of a rectangle ABCD of sides 8 cm and 6 cm are charges of +50μC, -30μC, +20μC 

and -40μC, respectively. AB is longer side and BC is the shorter side. Find the force (in N) 

on a +15μC charge placed at the center of the rectangle


1
Expert's answer
2021-04-05T11:10:12-0400


Let "q_1 = +50\mu C, q_2 = -30\mu C, q_3 = +20\mu C, q_4 = -40\mu C, q = +15\mu C", and "AB = 8cm = 0.08m, BC = 6cm = 0.06m".

Then, according to the Pyphagorean theorem:


"AO = BO = CO = DO = \sqrt{\left( \dfrac{AB}{2} \right)^2 + \left( \dfrac{AC}{2} \right)^2} = \sqrt{0.04^2 + 0.03^2} = 0.05m"

Then, according to the Coulomb's law, the force on "q" from "q_1" is:


"F_1 =k \dfrac{|q||q_1|}{AO^2}"

where "k\approx 9\times 10^9N\cdot m^2/C^2" is the Coulomb's constant.

Thus, find:


"F_1 = 9\times 10^9\cdot \dfrac{15\times 10^{-6}\cdot 50\times 10^{-6}}{0.05^2} = 2.7\times 10^{3}N"

The force on "q" from "q_2" is:


"F_2 =k \dfrac{|q||q_2|}{BO^2} = 1.62\times 10^3N"

The force on "q" from "q_3" is:


"F_3 =k \dfrac{|q||q_3|}{BO^2} = 1.08\times 10^3N"

The force on "q" from "q_4" is:


"F_4 =k \dfrac{|q||q_3|}{BO^2} = 2.16\times 10^3N"

The directions of these forces are show in figure above (like charges repel each other; unlike charges attract).

Adding the forces along "AD" and "BC" obtain:


"F_{AD} = F_1-F_3 = 1.62\times 10^{3}N""F_{BC} = F_4-F_2 = 0.54\times 10^{3}N"

These forces and their sum "F" are shown in figure below



According to the cosine law, the length of vector "F" is:


"F = \sqrt{F_{BC}^2 + F_{AD}^2 - 2F_{BC}F_{AD}\cos\angle BOD}"

The angle "\angle BOD" is:


"\angle BOD = 2\arccos\left( \dfrac{OE}{OB} \right) = 2\arccos\left( \dfrac{0.04}{0.05} \right) = 2\arccos0.8"

Substituting these values, obtain:


"F = \sqrt{(0.54\times 10^3)^2 + (1.62\times 10^3)^2 - 2\cdot 0.54\cdot 1.62\times 10^6\cdot 0.28}\approx 1.56\times 10^3N"

Answer. "1.56\times 10^3N".